Theory and Applications of Boundary Value Problems for ODEs

Douglas B. Meade

Department of Mathematics University of South Carolina Columbia, SC 29208

Phone:  (803) 777-6183

E-mail:  meade@math.sc.edu
WWW URL: http://www.math.sc.edu/˜ meade/

17 October 1997